OpenStudy (firejay5):

Show work & explain steps! :D Medal will be given! 7.

OpenStudy (firejay5):

\[\frac{ 2x + 1 }{ 3 } - \frac{ x - 5 }{ 4 } = \frac{ 1 }{ 2 }\]

OpenStudy (ajprincess):

First we need to find a common denominator for the fractions. To do that we have to find the l.c.m of 3, 4 nd 2. what do u think is the l.c.m? @Firejay5

OpenStudy (firejay5):

12

OpenStudy (ajprincess):

ya. good.:) Nw what do you get when u multiply each fractions by 12.

OpenStudy (firejay5):

the denominators cancel out

OpenStudy (ajprincess):

what remains after they cancel?

OpenStudy (firejay5):

the numerator

OpenStudy (firejay5):

8x + 4 - 3x - 15 = 6

OpenStudy (ajprincess):

is it -15?

OpenStudy (firejay5):

+ 15

OpenStudy (ajprincess):

ya:)

OpenStudy (firejay5):

Does the sign change when your multiplying/dividing or subtracting/adding???

OpenStudy (ajprincess):

it is \[-12*\frac{ x-5 }{ 4 }=-3(x-5)=-3*x-3*(-5)=-3x+15\] Does this help?

OpenStudy (firejay5):

so what ever sign